OpenStudy (theviper):

The radii of two circles are \(19cm\) & \(9cm\) respectively. Find the radius of the circle which has circumference equal to the sum of the circumferences of two circles.

OpenStudy (theviper):

@ash2326 help plz :)

OpenStudy (theviper):

@experimentX plz help :)

OpenStudy (experimentx):

hmm ... begin with Let the radius of circle be r, then circumference is ??

OpenStudy (theviper):

2\(\pi\)r

OpenStudy (theviper):

?

OpenStudy (experimentx):

yes ... and find the circumference of those two circles.

OpenStudy (theviper):

Bigger = 2\(\pi r\) = 2*22/7*19 = 836. Smaller = 2\(\pi r\) = 2*22/7*9 = 396.

OpenStudy (experimentx):

yes ... now let's add it up. 2 pi r = 2 pi r1 + 2 pi r2 <-- if you hadn't changed ... it would have been better.

OpenStudy (theviper):

oh but i had done a mistake Bigger = 836/7 Smaller = 396/7

OpenStudy (experimentx):

no .. don't change pi into 22/7 ... just keep it symbolic.

OpenStudy (theviper):

oh k

OpenStudy (theviper):

38\(\pi\)+18\(\pi\) = 56\(\pi\)

OpenStudy (experimentx):

yes ... euquate it to, 2pi r = 56 pi

OpenStudy (theviper):

r=56/2=28

OpenStudy (theviper):

now wht ?

OpenStudy (experimentx):

hmm .. it seems you are right. \[ r = r_1 + r_2 \]

OpenStudy (theviper):

Oh i got my answer Thanx Bhaiya :D

OpenStudy (experimentx):

hmm ... you are welcome :D

OpenStudy (theviper):

going to ask another one :)